WEIGHTED AVERAGE METHOD: Hurst Companys beginning inventory and purchases during the fiscal year ended December 31, 20X2 were as follows Date Event Units Unit Price Total Cost January 1, 20X2 January 12 February 28 June 29 August 31 October 29 November 30 December 21 bea. Invento 1st purchase 2nd purchase 3rd purchase 4th purhcase 5th purchase 6th purchase 7th purchase 1,50010.00$ 15,000.00 500$ 11.50 5,750.00 600 14.50$ 8,700.00 1,200$15.00 18,000.00 800$16.50$ 13,200.00 300$18.00$ 5,400.00 700$ 18.50 S12,950.00 400 20.00 8,000.00 $87,000.00 6,000 There are 1,200 units of inventory on hand at December 31, 20X2 Required 1) Calculate the total amount to be assigned to the Cost of Goods Sold for the year ending 20X2 and the ending inventory on December 31 under each of the following periodic inventory methods: a) Weighted Average Method (round calculations to 2 decimal places) 1c) Weighted Average Method Cost of Goods Sold x Unit Price E Cost of Ending Inventor Date Event Total Unitsx x Unit Price Total January 1, 20X2|Beg. Inventory
Answer #1

Weighted average cost per unit = $87000/6000 = $14.5

Cost of goods sold
Units × unit price = total
4800 × 14.5 = $69600

Units sold = begining inventory + purchase - ending inventory

= 6000 - 1200 = 4800 units

Cost of ending inventory
Units × unit price = Total
1200 × 14.5 = $17400

I habe given complete solution to your problem, in case of any query please comment and

PLEASE DO GIVE POSITIVE RATING